### v4

# Phaedrus: Predicting Dynamic Application Behavior with Lightweight Generative Models and LLMs

Abstract.
Application profiling is an indispensable technique for many software development tasks, such as code and memory layout optimizations, where optimization decisions are tailored to specific program profiles. Unfortunately, modern application codebases exhibit highly variant behavior across different inputs, creating challenges for conventional profiling approaches that rely on a single representative execution instance. In this paper, we propose Phaedrus, a new compiler-assisted deep learning framework designed to predict dynamic program behaviors across varied execution instances, specifically focusing on dynamic function call prediction.Such predicted call sequences are then used for producing optimized code pertinent to a given input.
Traditional profile-guided optimization methods struggle with the input-dependent variability of modern applications, where profiling on different inputs yields divergent application behaviors. To address this, Phaedrus proposes two new approaches: Application Behavior Synthesis, a profile-less approach where Large Language Models (LLMs) directly infer dynamic functions based on source code & static compiler analysis, bypassing the need for traditional profiling, and Application Profile Generalization, which uses generative models trained on compressed and augmented Whole Program Path (WPP) based function profiles to predict application behavior under unseen inputs. Our experiments show that Phaedrus can achieve upto $10^{7}X$ reduction in WPP function profile sizes, can predict most frequently executed functions that cover upto 85-99% of the execution time, along with an average of 13.19% (upto 65%) reduction in application binary size, and an average of 6.08% (upto 20%) performance improvement over the traditional profile-guided optimization, without any execution.
copyright: acmlicensedisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Application profiling and execution tracing are crucial in the software development process. Popular performance enhancement techniques such as profile-guided optimization (PGO) (radigan1995integer; grove1995profile; gupta1997resource) involve generating execution traces for applications and then performing optimization decisions specific to the collected profile, such as code transformation (chang1991using; visochan2022method; sydow2020towards; joshi2004targeted), data layout restructuring and reorganization (pettis1990profile). Application profiles and traces are also used to accomplish a wide variety of tasks such as estimating resource usage for scheduling and resource allocation (agelastos2015toward; song2009energy; viswanathan2011energy; rashti2015wattprof; agelastos2016continuous; kestor2013enabling), memory management for scratchpads, caches, and other memory hierarchies (cho2009adaptive; ansari2009profiling; bruno2017polm2; agarwal2017thermostat; xu2024prompt), memory simulation (uhlig1997trace; uhlig2001trace; rico2011trace; hassan2007synthetic) and debugging (arnold2007stack; nagarajan2012system; ning2017ninja; ning2018hardware; verma2020interactive). Such profile-driven approaches have been successfully utilized and deployed for over two decades.
Most profile-driven tasks rely on the users to manually select profiles representative of the application’s intended execution scenario. This ensures the usefulness of the profile analysis and optimization decisions during the application’s runtime across different inputs. However, recent work (chatterjee2022cas; mururu2023beacons; ahmed2017leveraging) has shown that most modern applications are complex and exhibit a high amount of input dependence, resulting in variable application behavior across different execution instances on various input data sets. This variant behavior is exhibited in all program components, including program branches, loops, and function calls. As a result, the application profile collected during a particular execution instance on a given input data set can be entirely different from another execution instance collected using a different input data set. A naive approach to tackle this problem would be to profile the application on multiple inputs - in practice, however, this ‘excessive’ profiling approach does not scale for workloads whose input domain is not numerable, such as graph applications. In addition, profile mixing can also be problematic for specific path-based optimizations that optimize along hot trace paths, pushing overheads to off-trace paths.
In this work, we study the problem of generating dynamic application behavior across different execution scenarios pertinent to different input data sets by capturing traces of application’s function calls and control flow behavior. Specifically, we explore the problem of hot function call prediction, where the objective is to obtain the set of computationally intensive (hot) functions called by the application during its runtime, specific to a given program input. In this paper, we propose Phaedrus, a new compiler-assisted deep learning framework which employs two different approaches to predict this dynamic program behavior: (1) application behavior synthesis, where a compiler-assisted LLM framework reasons about the applications’ algorithmic properties and extrapolates the application’s dynamic set of hot functions without any profiles; (2) application profile generalization, where lightweight generative models are synthesized to predict the application’s dynamic function call chains by learning a specialized representation of the smallest application profile. Phaedrus’s end goal is to use this predicted dynamic behavior to perform profile-guided optimization (PGO), such as changing the application function layout based on most executed functions to seek code size and performance enhancements.
Phaedrus Framework: In a nutshell, the idea of this work is to utilize generative models to predict the behavior of dynamic applications. The predictions from these models are then further used to perform profile-guided optimization. Towards this, the first new concept introduced in this work is application behavior synthesis, which attempts to reason about the dynamic program behavior without execution or collected profiles with the aid of Large-Language Models (LLMs). Owing to their remarkable successes in every domain, from natural language processing to software engineering, LLMs have been extensively used for generating and optimizing programs (ugare2024improving; gu2023llm; lin2024llm; liu2024exploring). However, the primary challenge that prevents the adoption of LLM in place of the mainstream compilation pipeline, is the absence of correctness guarantees (zhong2024chatgptreplacestackoverflowstudy; spiess2024calibrationcorrectnesslanguagemodels). Secondly, LLMs, being generative models at their core, cannot inherently perform static program analysis, which can enhance its decision-making capacities. Thus, naively querying the LLM for most code reasoning tasks fails spectacularly (as shown in fig. 1).
Figure 1. An example of unsuccessful dynamic program analysis with the LLM, given the source code and input files (520.omnetpp_r). The LLM is unable to pinpoint hot functions in the benchmark just by observing the source code & input files, and the output is quite unsatisfactory.
Compiler-Assisted LLM Reasoning: The first contribution of this work is the utilization of Large-Language Models, supplemented by the knowledge of static program properties obtained by compiler analysis. The Dynamis framework operates in three broad phases: first, filtering the wide corpus of LLM knowledge through domain knowledge deduction3.1); second, supplementing this understanding with program structural artifacts such as the universe of functions and call graphs (§3.2); and finally, fortifying the LLM’s reasoning with static compiler analysis3.3). While recent efforts have focused on using LLMs to generate optimized code (gao2024searchbasedllmscodeoptimization; qiu2024efficientllmgeneratedcoderigorous; rosas2024aioptimizecodecomparative), such code often requires extensive verification (liu2023; tang2024; ngassom2024). Unlike these approaches, Dynamis leverages LLMs only to identify input-specific profile-driven optimization opportunities, followed by traditional code generation techniques to elicit performance.
However, a downside of using LLMs for determining input-specific program behavior, is resource requirements needed for successful deployment and inference. For instance, deploying a state-of-art LLM such as LLaMA-2-70B with 70 billion parameters for inference, can take multiple NVIDIA A100 GPUs (zhou2024survey). Secondly, LLMs often have a fixed context window that determines the amount of tokens that it can consider to answer a certain query (8192 tokens for GPT-4). Large applications with hundreds of KLOC can easily exceed this limit, and lead to undesirable results. Most importantly, with the set of prompts used in the current work, we observe that LLMs are unable to reason about the dynamic sequence of function calls satisfactorily (§5.6). In the absence of accurate function call sequence, most path-based optimizations and execution forensics cannot be achieved.
To tackle these issues, we can use application profiles that capture the entire dynamic program behavior. This leads to the second important contribution of this work, application profile generalization, where the primary idea is to profile the application on a convenient single (small) input, transform it by adding program context (such as knowledge of possible program paths), and then train generative models on the compressed profile, which can then predict the application behavior on an unseen execution scenario (new input).
As a lightweight alternative to the LLM-based approach, in this work, we leverage Whole Program Paths (WPP) based function profile to make decisions based on the program’s runtime execution behavior. Whole Program Path Profiles (larus1999whole; ball2000using; zhang2001timestamped; tallam2005extended) captures the entire dynamic control flow of an application during its execution, and are superior in providing insight about the dynamic execution behavior than popular sampling-based trace collection methods (martonosi1993effectiveness; kessler1994comparison; novillo2014samplepgo; liu2016sample). However, the biggest challenge that prevents the adoption of WPP Profiles for profile-driven tasks in modern workloads is its intractability with respect to data size and volume (§5.5). Although compression of WPP profiles has been explored extensively in compiler literature (larus1999whole; zhang2001timestamped; burtscher2003compressing; zhang2002path; law2003whole; milenkovic2003stream; zhang2004whole; lin2007design), unfortunately such compression schemes can only be applied to a given WPP profile obtained from a single input data set at a time and cannot be used to get a generalized representation of profiles across different inputs (§2.3).
Lightweight Prediction of Dynamic Program Behavior4.1 - §4.4): An important contribution of this work is a new profile compression scheme, which is lossless and is consistent across different inputs. A WPP based function profile of the application is first generated on a convenient (smallest) input. While collecting the profile, function calls present inside loops are handled with smart instrumentation4.1) to avoid a high degree of redundant tokens. Compressible regions in the generated profile are identified by performing novel compiler analysis based on control-dependence and post-dominance relations (§4.2). These analyses are performed statically at the IR level to obtain compressible regions that are consistent across different inputs. The compressed trace is then augmented with additional possible control-flow paths and program context initially absent in the profile. This leads to a generalized application profile, which summarizes the program behavior across different execution settings.
The unified WPP based function profile serves as training data for this simple generative model, which, in turn, can predict program behavior specific to different inputs and execution phases. In this work, we show that these diverse execution behaviors can be learned with high prediction accuracy, by designing lightweight models that do not require any specialized hardware accelerators (GPU, TPU, etc.) for training / inference (§5.5). We also demonstrate the utility of predicting the dynamic function sequence (and the function set) called during execution in determining hot functions for optimizations (§5.5). The profile processing (generation, compression and augmentation), and the compiler analysis and model training, has been implemented as an end-to-end compiler-assisted deep-learning framework, Morpheus4).
The experimental evaluation (§5) of Phaedrus is done on the SPEC CPU 2017 benchmark suite (bucek2018spec). Our evaluation focuses on Dynamis and application behavior synthesis, we evaluate the prediction of hot functions and the coverage of several function calls (§5.1). We also show that Phaedrus can achieve application binary size reduction of upto 65% (13.68% on average) and performance improvement of upto 13.34% (2.8% on average). For Morpheus and application profile generalization, our evaluation focuses on the order of compression for WPP based function profiles (upto $10^{7}\times$ reduction), prediction of hot functions and its execution coverage and training overheads (§5.5). To the best of our knowledge, this is the first work to explore dynamic application behavior across different execution scenarios and leverages compiler-assisted lightweight deep learning models and LLMs for predicting new program behaviors for profile-guided tasks. .
2. Background and Motivation
In this section, we first describe the describe the rationale behind leveraging LLMs for predicting dynamic application behavior. We then motivate the need for a lightweight profile-driven approach, and explain the necessity of utilizing whole program path (WPP) profiles, and generative models for predicting dynamic application behavior. We also elaborate on the need of an input consistent compression scheme, and incorporating program context in WPP profiles.
2.1. Can LLMs reason about the dynamic program behavior?
Large Language Models (LLMs) exhibit remarkable capabilities in terms of code understanding (fang2024largelanguagemodelscode; nam2024usingllmhelpcode; wang2023codet5opencodelarge) and analysis (fang2024largelanguagemodelscode; ma2024lmsunderstandingcodesyntax), making them valuable tools for developers. One of their primary strengths lies in understanding code and describing behavior. LLMs can analyze and explain the logic of code snippets, breaking down complex functions into more comprehensible explanations for developers. They are also capable of predicting potential runtime behavior (chen2024reasoningruntimebehaviorprogram) based on static code analysis, such as identifying loops, conditionals, and function calls that may impact the execution flow. By simulating execution paths leveraging their neural models, LLMs can infer what might happen under certain conditions, offering valuable insight into code behavior without running it. Furthermore, LLMs excel in reasoning about high-level code behavior (ding2024reasoningplanningllm) by identifying concepts like state changes (e.g., variable updates or branching behavior) and evaluating algorithmic implications, such as performance considerations and potential impacts of code modifications. This makes them useful for predicting time complexity, spotting inefficiencies, and suggesting areas for optimization (yang2024large; zhou2024surveyefficientinferencelarge).
Figure 2. An example of successful dynamic program analysis with the LLM, given the source code (519.lbm_r) and input files. The LLM is successfully able to pinpoint hot functions in the benchmark for the particular input file, along with the exact number of function calls.
However, LLMs have limitations that stem from their inability to perform deep dynamic analysis. Because they do not execute code (xue2024; yang2024llm), they cannot analyze real-time state changes, such as tracking memory usage during execution or handling complex input/output scenarios dynamically. Their pre-trained nature also makes it challenging to reason about concurrency and parallelism, which involves understanding multi-threaded behavior, race conditions, or potential deadlocks. This type of analysis (sun2024llmruntimeerrorhandler; chen2024reasoningruntimebehaviorprogram; krishna2024codellmdevkitframeworkcontextualizingcode) requires direct code execution and monitoring of parallel processes, which LLMs cannot achieve. Furthermore, LLMs lack the capability to handle real-time data changes or interact with code based on live data, as they primarily operate on static representations and their training knowledge. This limitation prevents them from fully simulating execution environments or adapting to live input changes, highlighting an important gap compared to runtime analysis tools.
Although LLMs are inherently limited in performing extensive dynamic analysis, particularly in complex scenarios involving real-time state changes, concurrency, and live data interactions, our framework demonstrates a path to extend their utility through strategic assistance. By combining static analysis techniques with the logical reasoning (parmar2024; li2024relational) and domain knowledge that LLMs inherently possess, we can bridge this gap effectively.
2.2. Compiler-Assisted LLM Reasoning: Combining World Knowledge with Static Analysis
Static compiler analysis provides a depth of insight that LLMs relying solely on static inference from code often struggle to match. By meticulously examining the structural components of code, static analysis can prioritize and highlight important functions based on their position in the control flow hierarchy and their interactions across various program components. For example, identifying functions with self or mutually recursive calls or functions nested deeply within loops that span multiple function calls can indicate computational hotspots. This hierarchical understanding, which includes data on call frequencies, loop nesting levels, and data dependencies, offers a precise view of performance-critical regions within a codebase. Such insights are valuable for understanding which functions might contribute the most to the overall execution time of an application.
When these compiler generated static artifacts are paired with an LLM’s capabilities in reasoning (sun2024llmruntimeerrorhandler; chen2024reasoningruntimebehaviorprogram; krishna2024codellmdevkitframeworkcontextualizingcode) about control flow, algorithmic complexity, and domain-specific behavior, a comprehensive analysis of code behavior emerges. The LLM can use static analysis insights to simulate runtime behaviors, assess interactions, and predict the impact of changes more effectively. For instance, a call graph might reveal complex function dependencies, while the LLM can interpret how these dependencies influence runtime performance or suggest optimizations. However, performing such analysis within the LLM itself can be computationally expensive, shifting focus from estimating hot functions to more generalized codebase analysis. Since it is an interactive tool, LLMs typically budget a certain time interval for getting the response back to the user and within that budget different phases within the LLM must fit. To offload the burden of analysis and to utilize that time for some other inferencing phases, we assist the LLM with logical cues such as the set of all functions used within the benchmark, function call graphs, and some hotness indicators such as whether or not called within loops etc. This approach reduces the model’s search space, enabling more efficient and targeted inference by emphasizing functions that are statistically and structurally significant. This combination of static analysis and probabilistic guidance empowers the LLM to deliver more accurate and efficient code insights, bridging gaps in its inherent capabilities (hadi2024).
2.3. Alternative Approach: Profile Generalization & Generative Models
As an lightweight alternative to the inference costs and context length restrictions of LLM, we explore whole program paths as a means to reason about application behavior across different inputs. The Whole Program Path (WPP) based function profile represents the trace of complete control-flow path exhibited by an application during its execution. The program path, can be recorded either in terms of low-level program entities such as basic blocks and instructions, or in terms of aggregate artifacts such as branch outcomes. In this work, we collect the WPP based function profile for the application’s dynamic function calls, on a single input. Fig. 3 represents the overall function call and control-flow behavior for a simple benchmark (519.lbm_r) in SPEC2k17 suite. The corresponding WPP based function profile data is also depicted in the figure. This data can be interpreted as a stream of tokens, where each token represents a dynamic function call at a particular timestep in the program execution. For a given execution timestep, the function call token is dependent on the previous state(s) of the execution. For instance, in the given example (Fig. 3), function call token $7$ (init_channel) is invoked by the another function call $2$ (main_initialize) which occurred seven execution timesteps previously. Thus, in a given WPP based function profile sequence, the prediction of upcoming tokens requires historical information (previous execution timesteps), and involves handling of variable context length.
Figure 3. A simple outline of control-flow behaviour and function calls in 519.lbm_r (SPEC2k17). The WPP based function profile data on the small input is also depicted.
In order to utilize this WPP based function profile data for profile generalization and accomplishing profile-driven tasks, we train generative models on WPP based function profile data, which predicts the upcoming tokens (function calls) based on a prefix of tokens, and simultaneously adjusts to other execution scenarios (different inputs). Deep-learning generative models such as Recurrent Neural Networks (RNNs) can estimate the probability of the next token (at timestep $t$) in a sequence, as function of the previous token (at timestep t1t-1) and an internal short-term memory (which propagates information across the sequence). The utility of RNNs stems from the fact that the output token in the previous execution timestep acts as the input token for the current execution timestep, which simplifies the process of querying and inferencing the model. Popular generative models such as GANs (goodfellow2020generative), Diffusion Models (yang2023diffusion) etc, are extremely compute-heavy ($\sim 10^{20}$ FLOPS per epoch), and require extensive resources for training (100s of GPU hours). In contrast, we show that simple RNNs with one hidden layer (vanilla RNNs) are capable of learning the dynamic function call chains (§5.5).
However, there are two major challenges in directly adapting such an approach. The first issue is the scalability aspect of WPP based function profile traces - the disk size of function call WPP based function profile data ranges from a few GBs to $10^{2}$, which can lead upto billion of function call tokens (Table 21). For training RNN models, this data needs to be tokenized and encoded into tensors, which makes the dataset size unscalable (Fig. 23).
Need for an input-consistent compression scheme: In order to tackle this scalability issue of WPP based function profiles, it’s essential to compress and remove redundancies. A natural approach for compressing WPP based function profiles would be to simply encode frequent substrings of tokens into specific IDs, and then replace their occurrence in the profile accordingly. For example, in Fig. 3, the token substring $\{10\to 11\to 12\}$ inside the loop can be encoded as token $\{10^{\prime}\}$, reducing the size of WPP based function profile trace. Unfortunately, such substring-based strategies eliminates the control-flow semantics from the WPP based function profile during the training process. For instance, during the model’s testing, if the path inside the loop is $\{11\to 12\}$ (both if-statements false), for a different program input, then the compression encoding would not be useful, since substring $\{10\to 11\to 12\}$ is absent. Furthermore, the path encoding must be consistent across WPP based function profiles generated on different inputs, since the trained model needs to be tested on unseen, compressed WPP based function profiles. Thus, for successful compaction and generalization of WPP based function profiles, the compression scheme must not be dependent on a specific program input.
Incorporating Program Context: The second challenge arises from the fact that in the its current format, the WPP based function profilee does not include the full context of the program. For instance, in the example provided in Fig. 3, the conditional program path $\{8\to 8\}$ (if-statement false) is not reflected in the WPP based function profile trace generated on the small input (SPEC-test). In general, the possible program paths which are not expressed in a given WPP based function profile, needs to be provided in the training data for the generative model. Without the knowledge of program semantics and possible program paths, the raw WPP based function profile cannot be used directly for training generative models, and profile generalization.
Morpheus tackles the first challenge by devising a compression scheme that compacts tokens based on their control-flow behavior. It performs static branch analysis to determine tokens that can be chained together (compressible regions) for path encoding. In order to obtain input-consistent path encoding, Morpheus performs the path analysis directly on the source code. For instance, in the above example (Fig. 3), loops paths $\{8\}$, $\{11\to 12\}$ and $\{9\}$ would be given separate path encoding, because of the if-statements inside the loops. Morpheus also performs depth-first traversals on the function-call graph to handle cases with nested conditionals. Secondly, the lack of program context in the WPP based function profile is solved by augmenting possible program paths in the generated WPP based function profile. In the current example, sub-path $\{8\to 8\}$ is encoded and added in the compressed WPP based function profile to provide the complete context of the program to the generative models. The complete workflow of Morpheus, including its front-end which is responsible for compression & augmentation of WPP based function profile, and the back-end which trains the generative models, is presented in Fig. 4.
Figure 4. Morpheus Compilation Framework. The front-end of Morpheus is build as integrated set of LLVM compiler passes, while the back-end leverages deep learning framework Pytorch to data processing and training the generative model.
2.4. Basic Terminologies
In this section, we provide a formal definition of two classes ofhot functions, and some fundamental compiler terminology that is used extensively throughout the paper. Please see appendix A.
3. Dynamis: Using LLMs to Predict Program Behavior without Application Profiles
This section introduces Dynamis, compiler-assisted LLM framework which predicts the applications dynamic hot functions for a given input, without execution. The front-end of Dynamis consists of a set of LLVM-based compiler passes that generates the program artifacts required for LLM reasoning, while the back-end focuses on utilizing these predictions for performing superior profile-guided optimizations. In this paper, we are exploring Dynamis’ capabilities to determine two independent classes of hot functions: most frequently executed hot functions and runtime dominated hotspot functions.
Figure 5. Dynamis: A compiler-assisted LLM framework for predicting dynamic application behavior without any execution. The front-end of Dynamis provides the supplemental compiler analysis required by LLMs to reason about input-specific hot functions, while the backend involves leveraging those predictions to perform optimized compilation using LLVM’s profile-guided optimization (PGO) pipeline.
3.1. Phase 1: Filtering Search Space with Domain Knowledge Deduction
The first challenge in applying LLM-based reasoning across whole applications is navigating the overwhelming search space of thousands of functions, among which certain computation kernels dominate the application runtime. In the absence of additional guidance, it is difficult to distinguish between the actual kernels which are computationally intensive and the driver/utility functions which encapsulate the program hotspots. This makes it challenging for an LLM to isolate the true bottlenecks for optimization. Dynamis addresses this challenge by first deducing the domain of the given application (such as molecular dynamics, fluid mechanics, etc) with prompting. Once the domain is identified, algorithmic cost models and domain-specific principles are applied to predict which operations are likely to be the true bottlenecks. These deductions are then used to filter the effective set of candidate hotspot functions, ensuring that subsequent analysis focuses on the most computationally relevant regions of the code.
Our initial experiments showed that once the application domain is established, a direct correlation can be established between algorithm-specific cost models, and candidate hotspot functions. For instance, 510.parest_r510.parest\_r from SPEC2k17 (bucek2018spec) implements a finite element solver for biomedical imaging (Listing 1). Based on this domain knowledge, the LLM can immediately deduce that quadrature evaluations, gradient computations, and scatter operations are computationally intensive. This leads to the isolation of s_val, s_grad, and cpy_matrix as potential hotspots, sidestepping the high-level assemble_matrix wrapper. Similarly, 505.mcf_r505.mcf\_r from the same benchmark suite adopts a primal network simplex algorithm to solve the minimum cost flow problem (Listing 2). This deduction enlightened the LLM to attribute runtime to the domain specific constructs such as pricing (cost_compare), sorting (spec_qsort), and implicit-arc maintenance (bea_is_dual_infeasible) functions that dominate execution at scale, instead of the top-level driver primal_net_simplex. We observed a similar pattern in other application domains, such as molecular dynamics (neighbor list construction and electrostatics kernels), fluid mechanics (collide-and-stream sweeps), and image processing (convolution, and resampling), where domain knowledge consistently filtered out high-level drivers in favor of the computational kernels.
1void assemble_matrix() {
2 for (cell : mesh) {
3 for (q_point : quad_points) {
4 double phi = fe_values.s_val(.);
5 Tensor grad = fe_values.s_grad(.);
6 l_mat(i,j) += phi * grad;}
7 cpy_matrix(l_mat, g_mat);}}
Listing 1: Finite Element Solver code snippet from 510.parest_r in SPEC2017 (bucek2018spec). While the wrapper dominates inclusive time, the true hotspots are the inner quadrature and scatter kernels, based on domain knowledge.
1void primal_net_simplex() {
2 while (!optimal) {
3 primal_bea_mpp();
4 spec_qsort(.., cost_compare);
5 update_tree();
6 price_out_impl(); }}
Listing 2: Primal Simplex code snippet from solving minimum cost flow in 505.mcf_r in SPEC2017 (bucek2018spec). Despite enclosing the optimization loop, the runtime is dominated by the inner pricing and sorting routines
Furthermore, leveraging application domain knowledge allows Dynamis to apply algorithm-specific cost models, such as $O(nz)$ for sparse matrix–vector products, $O(N^{2})$ for neighbor-list generation, and $O(N\cdot r^{2})$ for convolutional operators, etc. This enables Dynamis to reason about scaling behavior and identify algorithmic bottlenecks across a wide range of applications. Note that the domain itself is inferred from several programmatic constructs, such as file structures, function signatures, symbolic constants, algorithmic patterns, and control-flow context. To validate this, we performed experiments with obfuscated function names and removed comments - the LLM still managed to infer the right domain by relying on algorithmic structure and control-flow regularities.
3.2. Phase 2: Supplementing with Program Semantic/Structural Artifacts
The second challenge in leveraging LLMs for arises when there is an ambiguity between the ‘moderately hot’ functions and dormant functions, which do not contribute to the program execution. Optimizing an application often requires a fine-grained view of program’s execution - from the dominant hotspots to secondary routines that still contribute meaningfully to runtime. The inclusion of domain knowledge narrows the space the to plausible kernels, but it can lead to overemphasizing a small set of domain-specific kernels or overlooking moderately expensive functions that accumulate cost across iterations. To address this, Dynamis augments its analysis with semantic and structural artifacts derived from static analysis, which includes complete function list, the program call graph, and the set of recursive functions with their transitive closures.
1void ResizeImage(Image *img) {
2 HorFilter(img); // medium hot
3 VertFilter(img); // medium hot
4 SyncCacheViewAutPix(); // cold
5}
Listing 3: Simplified Code snippet from imagick in SPEC2017 (bucek2018spec). The call-graph traversal from hot kernels allows Dynamis to isolate ‘moderately’ hot functions
The call-graph sheds light on caller–callee relationships, clarifying how the apparent cost of high-level drivers is distributed among their callees, and distinguishing moderately hot kernels that accumulate significant work from functions that remain largely dormant. This relationship is further refined with the inclusion of recursive functions, and helps the LLM to account for call multiplicity, which essentially identifies functions due to repeated invocation along multiple paths. For instance, Listing 3 shows a simplified code snippet from 538.imagick_r538.imagick\_r, where the call-graph highlights how filter kernels (HorFilter and VertFilter) and cache operation (SyncCacheViewAutPix) from the domain specific hotspot function ResizeImage. In this case, the both the filter operations are moderately hot for the large input, while the cache sync remains relatively insignificant. By traversing these caller–callee relationships, the call graph provides a graded view of functions—from dominant hotspots to moderately hot kernels to dormant routines, that better guides downstream optimization opportunities.
3.3. Phase 3: Fortifying LLM with Static Compiler Analysis
The addition of function call graphs and the recursive functions helps the LLM narrow down the function hotspots through caller–callee relationship, but does not quantify how often a kernel is actually exercised across different execution paths. For instance, two callee functions may appear equally central in the call-graph, but their effective runtime contributions can diverge sharply across different iteration of a loop nest. To address this issue, Dynamis augments the LLM with a static measure of interprocedural loop depth, which tracks how deeply a function is nested in loops both within its own body and across its transitive callers (definition LABEL:def:interloopiness). This provides an architecture-agnostic way to estimate repetition: functions invoked at deeper nesting levels or through recursive structures are more likely to dominate execution time.
Dynamis’ algorithm to estimate the inter-procedural loop depth is presented in the appendix (Algorithm A1). It computes the interprocedural loop nesting depth of functions by first recording intra-procedural nesting at each call site, then propagating these depths across the call graph using a longest-path calculation from the program’s main entry point. In practice, our experiments showed that combining all three phases puts Dynamis in the best position to determine computation hotspots of an application. For instance, in 519.lbm_r519.lbm\_r from SPEC CPU2017, a lattice–Boltzmann solver, the repeated full-grid sweeps consistently surfaced the collide–stream kernel as the dominant hotspot, with interprocedural loop depth highlighting its repeated execution across timesteps. By weighting call-graph edges with loop depth, Dynamis was able to filter out shallow wrappers and library calls, while consistently surfacing the kernels that dominate execution across specific input sets.
3.4. Phase 4: Performing Profile-Guided Optimization
Equipped with the supplemental materials from the earlier phases, semantic program artifacts (function lists, call graphs, recursion sets) and static analysis, Dynamis prompts the LLM to generate a refined set of top-k computationally hot functions for the application under a given input. As mentioned earlier, Dynamis generated two classes of independent predictions: Top-k frequently executed functions and Top-k runtime dominated hotspot functions. Both classes of predictions are required for contrasting the LLM’s predictive power, since they serve different optimization goals—one characterizes control-flow frequency, while the other identifies performance-critical bottlenecks.
After generating these predictions, Dynamis translates them into LLVM profdata files, which contains function hashes, basic block execution counts, branch probabilities in an indexed profile format. Dynamis searches for each hash corresponding to a predicted hot function, and then manipulates the counters based on the function ranking. This LLM-modified profdata is then passed onto LLVM’s profile-guided optimization framework, where it acts as a guideline for several optimization decisions such as function inlining, basic block placement (block layout), branch probability optimization, loop optimizations (unrolling, vectorization and peeling), function splitting (hot/cold splitting), register allocation and speculative optimization. Dynamisprofdata file ensures that the functions predicted by the LLM are boosted during PGO in the order of their predicted hotness.
4. Morpheus: Generalizing Whole Program Path Profiles
This section introduces the Morpheus Compiler Framework, which predicts the application dynamic function calls by generalizing its whole program path profiles (Fig. 4). The front-end of Morpheus is a LLVM-based instrumentation framework that generates the WPP based function profile in a compacted format, and performs static analysis to identify compressible regions in the WPP profile, and determines the paths for incorporating program context. The compacted WPP based function profile is generated by instrumenting the program loops, and the performing various online compression strategies. This allows Morpheus to replace the redundant function call sequences, which in turn, reduces the bias in WPP based function profiles. The static analysis for determining compressible regions is performed at the intermediate-representation (IR) level in order to obtain a compression scheme $\{path:code\}$, which can be applied on profiles generated by any program input. We refer to this scheme as input-consistent compression scheme.
After applying the compression scheme, the unseen program paths are augmented to generate a compressed, unified version of the WPP based function profile. This unified WPP based function profile serves as the training data for the Morpheus’s back-end. The profile is encoded into tensors, and then used to train lightweight generative models. The trained generative model can be leveraged to predict the application’s behavior on unseen inputs.
4.1. Collecting Whole Program Profiles: Smart Loop Instrumentation
The primary reason behind the intractability of WPP based function profiles in terms of their size, is the presence of highly redundant tokens and sub-paths. These redundancies are the result of the iterative behavior (deeply nested loops) of the program during its execution. Such data redundancies create extreme imbalances in the token distribution, which further causes significant bias in the WPP based function profile. This inhibits the downstream task of training generative models on the WPP based function profile. Morpheus tackles this problem by performing smart-loop-instrumentation and online compaction, during the generation of WPP based function profiles. The intuition here is that instead of directly logging function tokens in the WPP based function profile during nested loop executions, they are first compacted and summarized into lossless representations, while preserving their relative order in the program context.
Morpheus implements the smart-loop-instrumentation scheme by first instrumenting the loop headers, and the loops exit blocks, which trigger the activation and de-activation of a loop phase, with the help of special flags. Any function call tokens produced during loop phase are then stored into a temporary buffer. The contents of this buffer are subjected to online compaction whenever all loop phases in the application are deactivated. The online compaction scheme is described in Appendix (Algorithm 2). The distribution of the temporary buffers is first analyzed to see if it has been observed previously (lines 3 - 6). For compacting the buffer contents, the longest, repeating, unique and non-overlapping function call sequence is obtained, by designing a dynamic programming-based approach. This call sequence is then removed from buffer, and a special priority number is attached to the call sequence . The purpose of this counter is to determine the relative order of the longest call sequences. The algorithm is repeatedly invoked in the temporary buffer, until the content of the buffer are empty. Each longest, repeating call sequence is then logged in the WPP based function profile (in their respective order), with a begin and end markers to indicate the loop phase.
4.2. Identifying Compressible Regions: Static Program Analysis
The compressible regions in the WPP based function profile represents a sequence of function calls that can be grouped together into a single call. In absence of any control-flow (program branch), which can lead to variant program behavior, the entire sequence of function call tokens can be replaced with a single token. A naive approach to achieve this would be to simply group tokens which are not affected by any program branch. However, such an approach would miss compression opportunities for tokens which are nested inside branches. For determining the compressible regions in the WPP based function profile, Morpheus leverages the following insight regarding the function calls and branch targets, which can be stated and proved as a lemma:
Lemma 4.1.
In the absence of inter-procedural branches, function calls that are control-dependent on the same branch, and are present on the path of same branch target, can be compressed into a single token.
Proof.
Please see appendix C1. ∎
In order to determine control-dependence relationships, Morpheus first analyzes all function calls, branches (including switches) in the source code, and constructs the post-dominator tree which is used for detecting control dependence. The control dependence of each function call is checked with both the nearest branch and switch statements.
For simultaneously computing the control-dependence between function calls and branches, and identifying the specific branch targets on which the control-dependence property holds, Morpheus implements a bit-vector based detection mechanism (Algorithm 4). The algorithm treats each branch target as an element in a bit-vector (line 5). If a function call post-dominates the $i^{th}$ branch target, then position $i$ of the bit vector is set to 1 (lines 10 - 11). Thus, if a function call post-dominates all (or none) targets of a branch, then the bitwise exclusive-OR (lines 15 - 20) of the bit-vector would be 0. In that case, the function call is not control-dependent on the branch (lines 24 - 26). Otherwise, the function call dominates atleast one branch target (but not all), and is control-dependent on the branch (lines 21- 23). The output of Algorithm 4 is a mapping between every function call in the program and the corresponding branch target on which they are control dependent on (control independence is denoted as 0).
4.3. Input-Consistent Compression
The compressible regions in WPP based function profile profile can be identified by observing control-dependent function calls and the corresponding branch targets which they post-dominate. However, lemma 4.1 assumes the absence of any control-flow branches, which can potentially be present in callee functions, called by control-dependent functions. In order to reason about interprocedural branches for compression, Morpheus performs a modified depth-first traversal of functions, starting from the main function. This process is highlighted in Algorithm (Appendix C). For every intermediate callee function called by the caller function, this traversal takes into account the branch targets of caller functions, and checks if the callee is further dependent on some other branch. This processes is repeated recursively for all caller-callee functions (lines 8-9). After handling inter-procedural branches, Morpheus then checks the regions of the call stack which are dependent on the same branch targets, and encloses them into a single path.
The path encoding are then matched with sub-paths in the given WPP based function profile, where each instances of the sub-path are replaced with the corresponding path code. Our experiments have shown that in practice, this does not affect the compression drastically. The input-consistency of this compression scheme results from the establishing the path encoding directly from the application intermediate representation, instead of analyzing individual program profiles.
4.4. Augmenting Whole Program Profiles
The final step in the creation of an unified, compressed profile is the inclusion of program context in compressed profile. A single WPP based function profile represents the control-flow paths taken by the program for a particular input. It inherently lacks the knowledge of program branches, and conditionals, and the program context, which can lead to diverse program behavior. Morpheus tackles this issue by performing path analysis, and then augmenting the unseen paths in the compressed profile.
Figure 6. Simple example of augmenting unseen program paths in the WPP based function profile. The paths containing the set of unseen function $\{F_{4},F_{5}\}$, are obtained from traversing the call-graph starting from each unseen node. The traversal stops at the first seen node.
To achieve this, Morpheus first statically obtains a list of all the functions which are present in the application. For a given WPP based function profile, it then computes the set difference between the static function list and the observed functions in the WPP based function profile. For each (unseen) function in the set difference, it then performs a depth-first traversal of the function callgraph. The traversal terminates at the first function that is already present in the WPP based function profile. This operation results in a chain of unseen functions, which are then appended at the end of the compressed WPP based function profile. Fig. 6 depicts a single example where the WPP based function profile is augmented with the unseen program paths. Each augmented path is appended with start and end marker tokens, which preserves the context of the original trace. These marker tokens are also intended for indicating the generative model to adapt to input-specific path behavior.
4.5. Deducing Dynamic Function Calls: Sequence Prediction Problem
Dynamic function call prediction involves computing the sequence of function calls during a program execution. Specifically, the goal is to develop a prediction mechanism that can determine the function call $F_{t+1}$ at a timestep $t+1$, based on the program’s program execution state $h_{t}$, at timestep $t$. This can be expressed as sequence prediction problem with learnable weights $\mathbb{W}$, and bias $\mathbb{B}$:
(1) $\underbrace{F_{t+1}}_{\text{next call}}=\mathbb{W}\underbrace{h_{t}}_{\text{program state}}+~~\mathbb{B}$
For predicting the program’s dynamic function calls, the program’s execution state at timestep $t+1$, can be captured as a combination of function call $F_{t}$ at previous timestep $t$, and the program’s historical execution context (function calls $F_{0},F_{1},...,F_{t-1}$). The program state $h_{t+1}$ can then be expressed the following recurrence relation $f_{W}(h_{t},F_{t})$:
(2) $h_{t+1}=\underbrace{f_{W}}_{\text{activation function}}(h_{t},~~F_{t})$
In Deep Learning theory, the function $f_{W}$, which models the next program state based on previous state and function call, is known as activation function. Morpheus uses a rectified linear-unit (ReLU) as its activation function. The next function call token $F_{t+1}$, and the program state $h_{t+1}$ at timestep $t+1$ can thus be estimated with the help of Recurrent Neural Network (RNN). A typical RNN consists of neurons which can be distributed into input layer, output layer, and hidden layers. The function call tokens are first encoded into vectors, and passed on to the input layer. The hidden layer is responsible for modelling the program state by storing relevant information from previous timesteps. In this work, Morpheus uses RNNs with a single hidden layer, which is one of simplest neural architecture in theory (vanilla RNN).
The utility of RNN also stems from the fact that can be adapted to predict function call sequences in a variety of scenarios. For instance, in case of one-to-one prediction scenarios, the RNN accepts a new token at a single timestep, and outputs a single token. For next timestep, this output token can be re-used as an input token, which can give it one-to-many flavor, meaning that with multiple tokens could be obtained from RNN with a single input token. For function call predictions, the single input token would be the token corresponding main function of the application.
5. Evaluation
The evaluation of Phaedrus is two-fold. The first set of experiments are designed to evaluate Dynamis’s application behavior synthesis, while the second set of experiments evaluate the application profile generalization approach by Morpheus. Specifically, we evaluate Phaedrus to answer the following set of research questions:
  • Can Dynamis’s LLM-based generative model predict most frequently executed dynamic functions for given inputs? Can Dynamis be extended to predict the runtime dominated hotspot functions? How do these predictions impact performance and code size?
  • What sources of additional information should be provided to Dynamis for achieving close-to-ground truth results? Finally, what fraction of function calls do the predicted functions cover compared to the ground truth when provided with additional inputs?
  • How effective is Morpheus’s smart-loop instrumentation mechanism in compressing the WPP based function profiles? Is it able to reduce redundancies in a lossless manner?
  • Can Morpheus’s RNN-based generative model predict most frequently executed dynamic functions for unseen inputs? How “lightweight” is the entire approach, in terms of training time and model size?
  • How much performance improvement & binary size reduction can be achieved by profile-guided optimization by using function hot sets generated by Morpheus & Dynamis? What are the tradeoffs between these two approaches?
Experimental Setup: The experiments were conducted on a Dell 7820 Precision Tower with Intel Xeon Gold 6230R processors. The Intel system has 2 sockets (13 cores each), 36 MB L3-cache, running Ubuntu 20.04.6 LTS. The compiler passes and algorithms in front-end of Morpheus & Dynamis were implemented in LLVM 15.0.0, while Pytorch version 2.4.0, was used in Morpheus’s deep-learning back-end with Python 3.8.10. GPT-5, Claude-4-Sonnet and GPT 4.0 were used as the LLM in Dynamis. The LLMs were invoked in an interactive environment by Cursor (cursor_ai_2024) .
Hyper-Parameters & Baselines: The RNN-based model used in Morpheus consists of single-layer, with 1000 hidden units. The training batch size was set to 4, 32, and 64 for benchmarks with $\leq 50,\leq 1000,>1000$ tokens, respectively. The learning rate was set to be $0.001$, dropout $0.2$, and each RNN was trained for 20 epochs.
PGO Baseline: To compare the PGO benefits of Phaedrus, we use two different baselines (1) O3-baseline: applications compiled normally with O3-flag (2) traditional PGO, where an application is first profiled on representative input(s), and then compiler transformations are performed based on the given profile. For this baseline, we profile the application on small input and use that profile to perform PGO.
5.1. Evaluating Dynamis: Application Behavior Synthesis
Our experiment of extrapolating the application domain by Dynamis showed a very high (almost perfect) success. With GPT 4.0, we observed that except 510.parest_r, the deduction is very close to the actual domain. However, when this experiment was repeated with GPT-5 and Claude-4-Sonnet, we observe that the domain deduction is always correct. This makes way for accurate reasoning of the algorithmic behavior of domain-specific applications. Table 1 in appendix shows the complete extrapolation of different domains by GPT 4.0-equipped Dynamis in Phase 1 (§3.1) across different benchmarks.
The effects of prompting the LLM to first infer domain knowledge is highlighted in Fig. 8. The most notable improvement was in benchmark 531.deepsjeng_r improvement, which immediately led to the identification of computationally intensive functions.
[Uncaptioned image] Figure 7. Dynamis’s prediction of top-15 most frequently executed functions in the three phases. The trend how Dynamis refines it’s predictions, once equipped with finer granularity of compiler knowledge. [Uncaptioned image] Figure 8. Percentage of most frequently executed function calls predicted by Dynamis, and compared with the actual execution (ground truth).
The predictions of most frequently executed hot functions improved significantly in phase 2 (§3.2), where the LLM was provided with program semantic artifacts. Specifically, providing function universe improved the results by 8.33%, notably in benchmarks 544.nab_r, 508.namd_r by 43.33%, while call graph, recursive functions and transitive closures enhanced the output by 2.8%, 5.56%, and 12.22% respectively. The function universe reduces the search space, while call graphs exposes caller-callee relationships, aiding in statically tracing control flow. We observed that recursive functions also lead to high-execution hotspots and transitive closures uncover long-range dependencies. These elements enhanced Dynamis’ ability to prioritize and predict computationally intensive functions.
Finally, in phase 3 (§3.3), we observe inter-procedural loop depth, further refines LLM estimates by indicating function hotness based on computational complexity. However, estimating the size of loop bound to a particular depth remains challenging. Although Dynamis works well on most benchmarks, limitations persist for 520.omnetpp_r and 525.x264_r, which lack correlations between static analysis cues and runtime behavior, severely making it harder for LLM to determine the most frequently executed functions.
5.2. Expanding Dynamis: Exploring Modularity with new LLMs
In order to investigate Dynamis’ performance with different LLMs, we repeated the experiments with two current state-of-the art ”thinking” LLMs: claude-4-sonnet and GPT-5. In these experiments, we designed the prompt to account for the ”functions where the applications spend the most amount of execution time”. This is different from our previous experiments, since we are shifting the focus from the set of most frequently executed hot functions to runtime dominated hotspot functions. Predicting runtime-dominated hotspot functions is a harder problem because execution time depends on complex interactions (such as memory patterns, data dependencies) rather than function invocation counts.
Fig. 9 shows Dynamis’ results on predicting the runtime-dominated hotspot functions, and Fig. 11 illustrates the percentage of application execution runtime covered by these hotspot functions. In our experiments, we considered any function accounting to more than 5% of the total application runtime as a ‘hotspot’ function. As observed, Dynamis’ hotspot prediction accuracies and coverage can broadly be divided into three categories:
Figure 9. Dynamis’ prediction of hotspot functions in the SPEC2017 suite for REF input. Functions that amount for more than 5% of the total application execution time are considered here. With the aid of domain knowledge and static compiler analysis, Dynamis is able to infer the computationally intensive hotspots for large applications without profiling or executing.
$\blacksquare$ Near Perfect Coverage ($80\%$ execution predicted): In 6 out of 13 benchmarks from SPEC2017, GPT5-equipped dynamis was able to extrapolate almost all of the runtime program hotspot functions. These are majorly compute-bound physical/chemical simulation benchmarks, where there exists a direct correlation between domain knowledge and the algorithmic complexity. For instance, in 508.namd_r508.namd\_r, the LLM first identified the application domain as molecular dynamics, and based on that, it knew that “atoms interact through long-ranged forces, which involves atom-pairs”, leading to $O(n^{2})$ scaling. It was further reinforced with the high inter-procedural loop depth ($\sim 214$) for the aforementioned computation. Similarly, in 519.lbm_r519.lbm\_r, the LLM extrapolated the algorithmic complexity $O(N\cdot T)$, after inferring that the application implements the Lattice-Boltzmann Method from fluid mechanics. The rest of the hotspots were then obtained by mapping the call graph to the corresponding functions, and incorporating the loop depths to highlight potential bottlenecks. Also, note that benchmarks from this category range from a few KLOC to $\sim 250$ KLOC.
[Uncaptioned image] Figure 10. Percentage of application execution runtime covered by Dynamis’s hotspot predictions for SPEC2017’s REF input. The highest coverage is obtained on compute-intensive applications, while memory-bound applications have lower prediction accuracy. [Uncaptioned image] Figure 11. Percentage of context window used by Dynamis for predicting hotspot functions. A state of the art LLM such as GPT5 barely requires a third of its context window to predict input-specific dynamic behavior.
$\blacksquare$ Moderate Coverage ($50-80\%$): 4 benchmarks (500.mcf_r500.mcf\_r, 510.parest_r510.parest\_r, 520.omnetpp_r520.omnetpp\_r, 541.leela_r541.leela\_r) exhibited an execution coverage of 60% on average. These benchmarks are applications where the hotspot functions are not confined to the primary algorithmic kernels but also include auxiliary and system-level routines that dominate at scale. For instance, in 505.mcf_r505.mcf\_r, the domain knowledge combined with the semantic program artifacts correctly identified 4/5 hotspot functions by zeroing in on the core algorithmic kernels. However, both the LLMs missed the utility sorting function spec_qsortspec\_qsort, which dominates at scale. Similarly, in 510.parest_r510.parest\_r, the LLM was able to identify algorithmic hotpots, but failed to reason about the various C++ template variants missing a couple of hotspots. One possible solution to improve coverage for such benchmarks would be to incorporate finer-grained static analyses—such as resolving function pointers and template instantiations.
$\blacksquare$ Minimal Coverage ($0-35\%$): The runtime hotspots in 523.xalanc_r523.xalanc\_r, are dominated by memory allocation/deallocation patterns, and object lifecycle management. This makes the application memory-bound, and the LLM’s domain expertise, which focused on determining algorithmic complexity, did not translate to actual hotspots. On the other hand, 500.perlbench_r500.perlbench\_r, exhibits a near-uniform distribution of execution time ($\sim 1-2\%$ each), with two outlier functions amounting to only a third of the runtime. Dynamis was able to infer the dominant outliers (S_regmatchS\_regmatch, perl_hv_commonperl\_hv\_common), leading to low coverage. 525.x264_r525.x264\_r has a similar distribution, with the longest function only lasting 4 secs.
5.3. Leveraging Phaedrus: Profile-Guided Optimization
$\star$ Runtime-Dominated Hotspot Functions: We also used Dynamis’ prediction of hotspot functions (§5.2) to perform PGO. The performance improvement over a normalized O3-baseline is shown in Fig. 13. Dynamis achieves a performance improvement of 6% and 3.63% over the traditional PGO baseline with claude-4-sonnet and GPT-5 respectively. An important point to note is that even though Dynamis’ predictions covered only 50–60% (Fig.9) of the total execution time (and less than 30% in some cases), its inferred hotspots were still broadly consistent with those observed in the ground-truth profdata profiles. This is particularly interesting because it suggests that even with runtime partial coverage, the framework captured the core computational patterns responsible for most of the runtime improvement.
To investigate this effect further, we analyzed the contents of the predicted profdata and the ground truth closely. Profiling tool gprof highlights expensive compute-bound routines that dominate wall-clock time, while profdata emphasizes fast but frequently executed functions. It became evident that functions dominating runtime are not always those most amenable to direct optimization. In some cases, performance can be improved indirectly—by transforming other functions that invoke these heavy kernels. For instance, in 520.omnetpp_r520.omnetpp\_r, the low-level function cMessageHeap::shiftup predicted by Dyanmis appeared trivial in gprof, yet it dominated execution counts in profdata. We observed that Dynamis’ reasoning consistently balanced between capturing both computationally intensive and optimization-relevant routines. In spite of being prompted to identify runtime-dominated hotspots, Dynamis still surfaced functions that were most likely to yield measurable performance improvement when optimized. This effect was also observed in the application binary sizes, where Dyanmis achieved an improvement of 5.19 over O3. This is achieved by transforming the function layout within the application binary. This optimization reduces the binary size by enabling superior function compression (hoag2024reordering), and can potentially improve performance as the new layout would minimize page faults by appropriately ordering them.
[Uncaptioned image] Figure 12. Dynamis’ hotspot predictions achieves an average performance improvement of 6.08% over traditional PGO, without any profiling. This is achieved by capturing both computationally intensive and optimization-relevant routines by Dynamis. The time shown is normalized with respect to O3. [Uncaptioned image] Figure 13. Dyanmis hotspot predictions achieves average codesize reduction of 4.53% & 3.26% over O3 with claude-4-sonnet and GPT-5 respectively. This is achieved by function re-layout transformation which enables superior function compression. The size shown is normalized with respect to O3.
$\star$ Frequently Executed Functions: We also leveraged Phaedrus’s outputs (both Morpheus & Dynamis) of most frequently executed functions to perform PGO. Fig. 17 & Fig. 15 shows the effect of performing PGO with Phaedrus’ outputs. As observed, we notice performance improvements comparable to traditional PGO (0.8% on average and upto 19.18% at max) on most benchmarks.
On taking a closer look into the performance trends, we find that on average Morpheus outperforms Dynamis’ frequently executed functions setting by 2% on average for both baselines. This is explained by the fact that Morpheus’s RNN model generates the unified trace in a timestep manner; rather than extrapolating the computationally hot functions, as Dynamis does. This highlights a fundamental difference in the methodology of both approaches: application behavior synthesis, a profile-less approach, is approximate than application profile generalization, which involves profile generation, compression and unification. This intuition is clearly reflected in Fig. 17 & 15.
Fig. 15 highlights the binary size reduction for Dynamis over both baseline. Dynamis outperforms the normal O3 baseline (no PGO) by 14% on average with up to 65% in 519.lbm_r519.lbm\_r. As mentioned earlier, the primary reason for size reduction is the high compression obtained by appropriate function layout by LLVM. Fig 17 depicts the binary size reduction over both baseline - in this case, we don’t notice any observable difference.
[Uncaptioned image] Figure 14. Dynamis’ frequently executed predictions achieves around 0.8% average improvement over O3. Being a legacy model, GPT 4.0 often missed the exact computational hotspots in large benchmarks. [Uncaptioned image] Figure 15. Dynamis’ frequently executed predictions reduced codesize by an average of 13.68% over O3, outperfoming the hotspot predictions.
[Uncaptioned image] Figure 16. Morpheus’s results for codesize reduction. An average of 17.63% improvement is observed, just as good as traditional PGO [Uncaptioned image] Figure 17. Morpheus’s performance improvements observed across different benchmarks. Baseline here is normal compilation without profiling. Morpheus achieves around 2.8% average improvement over the baseline.
5.4. Case Study: Real-World Benchmarks
In order to better demonstrate the capabilities of Dynamis on larger real world programs, we evaluated the framework on 8 popular graph processing problems ((beamer2017gapbenchmarksuite)) (BFS, CC, CC_SV, TC, PR, PR_SPMV, SSSP) with an undirected graph of degree 15 with 16777216 nodes and 268435177 undirected edges, and also with the popular compiler benchmark GCC, which has 1304 KLOC. Fig. 19 shows that Dynamis outperforms O3 baseline by 4.45% for GCC. It navigated the huge codebase by starting with the program inputs & flags, based on which it inferred the dominant compiler passes. Then it identified the ”known” heavy kernels within those passes (bitmaps and graph operations, points-to analysis, etc), which were then ranked by their expected algorithmic cost. This shows how Dynamis is able to adapt its reasoning for large codebase by tracing its thinking from the inputs.
[Uncaptioned image] Figure 18. Dynamis’ hotspot predictions on real-world programs showed that it can achieves a performance improvement of 4.45% and 9.09% over O3 and traditional PGO in GCC. The average improvement is -0.92% over O3, which is a slight slowdown. [Uncaptioned image] Figure 19. Dynamis’ hotspot predictions on real-world showed an average improvement of 6.95% over O3, and is slightly better than traditional PGO on average. The graph benchmarks showed the most reduction in size.
On the other hand, in the graph workloads, Dynamis switched back its default strategy of first leveraging domain knowledge (such as union-find and path compression in CC) to identify compute-heavy kernels, and then mapping this theoretical knowledge to the source codes. It also leveraged the call graph to filter low-level STL calls, then determined a ranking based on the given input. In this case, Dynamis was able to achieve an average reduction in the application codesize of 7.83% over O3, but it did not translate to significant performance improvement ($0.37\%$ over O3). A possible reason for this is that these programs BFS/PR/SSSP/CC are memory-bound and spend time on random memory access; reducing code size doesn’t impact bandwidth/latency.
5.5. Evaluating Morpheus: Application Profile Generalization
$\star$ Need for Compressing WPP based Function Profiles: Table 21 shows the total tokens and disk size for WPP based function profiles expressed in terms of application function calls for SPEC2k17 (bucek2018spec) benchmarks. As shown, the total function call tokens can range up to $10^{10}$ tokens per profile while incurring disk sizes of a few GB to $10^{2}$ GB. Without the compression techniques employed by Morpheus, such profiles cannot be analyzed. For instance, Fig. 23 summarizes the result of training a simple RNN model with one hidden layer on the uncompressed WPP data generated from the smallest program input, in SPEC2k17 Benchmarks. Out of 14 benchmarks, the training process on 7 benchmarks failed because of exceeding the tensor’s default memory capacity in the Pytorch framework, while 5 benchmarks had out-of-memory errors, and were terminated by the OS scheduler. Furthermore, Fig. 23 depicts the low ratio of unique tokens in WPP based function profiles. This indicates a significantly high rate of duplicate (and redundant) tokens, which can lead to excessive bias and unnecessary computations during training.
Benchmark WPP based function profile Disk Size Function Tokens Small Medium Large Small Medium Large 500.perl_r 305 MB 887 MB 8.2 GB 66.5 M 198.2 M 1.8 B 502.gcc_r 1.3 MB 11 GB 19 GB 248 K 2.2 B 4.3 B 505.mcf_r 2 GB 11 GB 58 GB 687 M 3.5 B 20.5 B 508.namd_r 25 MB 73 MB 1.6 GB 8.6 M 60.1 M 558 M 510.parest_r 110 MB 692 MB 38 GB 6 M 35 M 2.2 B 519.lbm_r 4 KB 4 KB 16 KB 75 919 6056 520.omnetpp_r 853 MB 14 GB 124 GB 4.4 M 43 M 378 M 523.xalanc_r 2.2 MB 854 MB 95 GB 421 K 127 M 18.9 B 525.x264_r 25 MB 25 MB 387 MB 3.8 M 3.8 M 40 M 526.blender_r 56 MB 8.8 GB 27 GB 10.6 M 1.8 B 5.3 B 531.deep_r 2.4 GB 20 GB 110 GB 510 M 5.1 B 39 B 538.imagick_r 248 KB 183 MB 90 GB 55 K 47 M 24 B 541.leela_r 2.3 GB 38 GB 169 GB 454 M 7.9 B 58 B 544.nab_r 4.7 MB 25 MB 2.6 GB 76 K 185 K 1.6 M 557.xz_r 252 KB 297 MB 2 GB 11.8 K 76 M 537 M Figure 20. Total Tokens and their corresponding Disk Sizes, for WPP based function profiles consisting of function calls in SPEC2k17 (K = $10^{3}$, M = $10^{6}$, B = $10^{9}$) [Uncaptioned image] Figure 21. Time Taken for Collecting WPP based function profiles for different Input Sizes (small, medium, large) in SPEC Suite. In most cases, WPP based function profile on a convenient input can be generated within 10 secs.
[Uncaptioned image] Figure 22. Results of training RNN models directly on the uncompressed WPP data in SPEC2k17 suite. The training was successful only on 2 benchmarks, while the rest (13 benchmarks) led to out of memory and memory allocation errors. [Uncaptioned image] Figure 23. Fraction of unique tokens in WPP data on all inputs in SPEC2k17. The plot indicates a very low fraction of unique tokens in the WPP based function profile, which is also inversely proportional to the input size.
$\star$ Smart-Loop Instrumentation: Table 25 depicts the size of WPP based function profiles collected after the smart-loop instrumentation scheme. As observed, the original WPP based function profiles that were originally in GBs, were reduced to mere KBs in most cases. Fig. 25 shows the ratio of reduction achieved in WPP based function profiles for all input sizes. For certain large profiles (531.deepsjeng_r, 538.imagick_r, 541.leela_r), the compression ratio is upto $10^{7}\times$, which points to the high degree of redundant tokens. This makes the WPP based function profiles scalable, and makes the way for further analysis for training generative models.
Benchmark Smart Loop Instrumented WPP based function profile Disk Size Small Medium Large 500.perl_r 16 KB 248 KB 600 KB 502.gcc_r 220 KB 2.3 MB 4.5 MB 505.mcf_r 4 KB 4 KB 4 KB 508.namd_r 4 KB 4 KB 4 KB 519.lbm_r 4 KB 4 KB 4 KB 520.omnet_r 21 MB 206 MB 1.8 GB 523.xalanc_r 116 KB 25 MB 4.9 GB 525.x264_r 8 KB 8 KB 20 KB 526.blender_r 260 KB 312 KB 320 KB 531.deep_r 8 KB 16 KB 16 KB 538.imag_r 8 KB 12 KB 16 KB 541.leela_r 12 KB 12 KB 16 KB 544.nab_r 4 KB 4 KB 4 KB 557.xz_r 4 KB 4 KB 4 KB Figure 24. WPP based function profile Disk Size after Morpheus’s Smart-Loop Instrumentation. In majority of the benchmarks, the profile sizes are reduced to KBs from GBs. [Uncaptioned image] Figure 25. Ratio of Reduction in Profile Size achieved by Smart-loop Instrumentation over Normal WPP based function profiles. The trend shows that Morpheus achieves higher degree of compression for largest input sizes, which is in line with the higher redundancies in larger profiles.
$\star$ Input Consistent-Compression & Augmentation: Fig. 27 depicts the results of the input-consistent compression scheme, where each compressible region is statically assigned an encoding. In majority of the benchmarks, the total regions are below 20. The largest number of compressible regions identified in a benchmark was 498 (526.blender_r). The average function call length of these regions was nearly 3 function calls per compressible regions across all benchmarks (Fig. 27).
[Uncaptioned image] Figure 26. Input-consistent compressible regions identified statically from SPEC benchmarks. [Uncaptioned image] Figure 27. Average length of input-consistent compressible paths for each benchmark.
[Uncaptioned image] Figure 28. Total tokens augmented in WPP based function profile for capturing program context. [Uncaptioned image] Figure 29. Final Token Count in WPP based function profile after input-consistent compression.
$\star$ Determining Computationally Frequent Functions: Fig. 31 shows the computationally hot functions predicted by Morpheus on unseen program inputs. As observed, the in all cases, Morpheus is able to predict close to 99% of the total function calls in the program, by only predicting a fraction of total functions present in the application. This is explained by the fact that the RNN-based model can effective capture the sequence of functions captured in loop-intensive regions, and is adapts to the unseen inputs by activating the tokens corresponding to the unseen augmented paths.
Fig. 31 shows the training time required by Morpheus to learn the generalized WPP based function profiles. As observed, Morpheus can complete the training process in minutes without deploying any specialized hardware. The fast training is a direct consequence of Morpheus pre-processing the WPP based function profile into input-consistent compressible regions, which reduces the total tokens. Furthermore, the redundancy removal by the smart loop instrumentation scheme reduces bias, thereby preventing the model results from being skewed towards certain functions.
[Uncaptioned image] Figure 30. Morpheus predicting of computationally most frequently executed functions, as fraction of total functions and execution covered. The trend shows that Morpheus is able to predict almost 99% of program’s execution with only a fraction of functions. [Uncaptioned image] Figure 31. Morpheus’s training time for generative models. The RNNs were trained for 20 epochs in each case, with varying batch sizes ranging from 4 to 64. Morpheus is able to complete training within minutes, without any specialized accelerators.
$\star$ Scaling of Morpheus with different application sizes: Fig. 32 shows Morpheus’s compressed profile generation time across applications of increasing size. As observed, the compression time grows log-linearly with code size. This is especially highlighted for large codebases such as 526.blender_r, where the major bottleneck is the online compaction scheme of detecting longest repeating call chains. Across all benchmarks, larger input datasets consistently take longer, confirming that Morpheus’s cost scales with both program and input size.
5.6. Dynamis Vs Morpheus: performance tradeoffs and considerations
In this subsection, we explore the utility and performance tradeoffs of both Dynamis and Morpheus; compared against one another. The primary necessity of a single profile-driven approach such as Morpheus arises from the fact that with the default set of prompts and model weights, the LLM is unable to provide any sequential ordering of the function calls during runtime. For instance, querying GPT 4.0 to reason about the dynamic call for the smallest input in 519.lbm_r generates the following trace (missing tokens represented in red):
$\text{LLM-generated Trace}:0,1,2,4,5,6,8,\{10,11,12\}\times 20$~~~~~
$\text{Actual Trace}:0,1,2,4,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4},5,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}5},6,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}6,7,7,9,}\{10,11,12\}\times 20,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}9,3,3}$
Figure 32. Morpheus’ Log-linear scaling of compression time with larger inputs and bigger codebases
For reference, the entire control flow is represented in Fig. 3. Thus, in order to obtain a clearer picture of an application’s dynamic sequence of function calls, we employ profile generalization. As described earlier in §5.5, Morpheus is able to achieve upto 99% of the program’s execution. However, achieving application profile generation and successfully leveraging it for profile-driven tasks still requires the users to generate a single application profile by executing the program on a representative input. Furthermore, there is an implicit assumption about the user’s necessary expertise to determine the convenient input set on which the application profile needs to be collected. For specific long-running applications, the task of program execution and data collection could potentially result in a bottleneck. For instance, Fig. 21 shows the distribution of time required to collect WPP traces. As observed in the figure, profiling certain execution instances can even take up to 3 hours (541.leela_r), which can complicate the process of profile generalization. Thus, we still need to employ profile-less approach such as Dynamis. Another noteworthy point in the previous example is that inspite of the obvious inaccuracies in the program trace, LLM was not only able to predict the hot functions (10, 11, 12), but also the frequencies accurately. We anticipate that with better prompting and sophisticated fine-tuning, LLMs can be used to bridge the gap in function call chain prediction as well.
6. Related works
Recent research in code optimization has transitioned from rule-based compiler heuristics to large language model (LLM)–driven code transformation and refactoring (gao2024searchbasedllmscodeoptimization; rosas2024aioptimizecodecomparative; ma2024llamocoinstructiontuninglarge; pomian2024furtherllmsidestatic; cordeiro2024empiricalstudycoderefactoring). LLM frameworks combine sequence generation with iterative refinement using evolutionary search, adaptive retrieval, and specialized prompting (liu2023promptingframeworkslargelanguage; wu2024evolutionarycomputationeralarge). Comparative studies show that LLMs can rival or surpass traditional compilers such as CETUS and PLUTO in optimization quality but often struggle with correctness and scalability on large real-world programs. Instruction-tuned models (ma2024llamocoinstructiontuninglarge) further address these issues through optimization-aware fine-tuning, improving reliability and consistency across diverse workloads.
In parallel, machine learning–based static profiling has emerged as an alternative to dynamic PGO, eliminating the need for costly runtime instrumentation and representative input collection. VESPA (vespa2021moreira) pioneered static frequency prediction for binary optimization in Meta’s BOLT (panchenko2018boltpracticalbinaryoptimizer), achieving up to 6% speedup over Clang –O3 by estimating basic-block frequencies directly from static features. GraaSP (vcugurovic2024graalsp) extended this concept to compiler integration within GraalVM (graalvm2020), embedding ML inference during compilation to remove the two-phase PGO cycle and delivering a 7.5% runtime gain. GraalNN (milikic2025graalnn) advanced the field further by employing Graph Neural Networks to model call-graph context, refining predictions across parsing and inlining phases for greater accuracy. Building on these insights, our work explores input-specific static profiling without execution, using LLMs to infer hot regions based on algorithmic structure, input size, and termination behavior. Unlike prior systems targeting speed in Java or binary infrastructures, our framework focuses on code optimization in C/C++ workloads, evaluated against PGO baselines on SPEC CPU 2017. Though not directly comparable, our results demonstrate that LLMs can approximate—and often surpass—traditional profiling accuracy while offering fine-grained, input-aware guidance.
On the other hand, the concept of capturing comprehensive dynamic program behavior through Whole Program Paths (WPP) has been extensively explored in the compiler literature (larus1999whole; zhang2001timestamped; burtscher2003compressing; zhang2002path; law2003whole; milenkovic2003stream; zhang2004whole; lin2007design). WPPs provide a single, compact description of a program’s entire dynamic control flow, including interprocedural calls and loop iteration, which is essential for finding ”hot subpaths” for optimization. While techniques for compressing WPP profiles are well-developed, they are inherently limited to analyzing a single profile instance generated from one specific input data set. Our approach, in contrast, attempts to generalize across multiple profiles by using a novel, lossless compression scheme that is consistent across different inputs, enabling the prediction of program behavior for unseen execution scenarios.
7. Conclusion
This paper introduced Phaedrus, a novel framework that advances profile-driven software optimization through predictive modeling of dynamic application behavior. Phaedrus offers a unique hybrid approach that combines Application Profile Generalization and Application Behavior Synthesis. Our framework uses generative deep learning models and large language models (LLMs) integrated with static compiler analysis to accurately predict function calls and control flow behavior, enabling adaptive optimizations without the need for exhaustive profiling across diverse inputs. We show that LLMs provide a good prediction capability of hot functions in an input specific manner that yield performance that is comparable to and slightly better than traditional PGO and code size reduction comparable to traditional PGO. RNN based framework improves performance over traditional PGO by 2.8% although an initial investment of doing profiling and model generation is needed. Due to its near perfect accuracy, a key use of Morpheus could be to perform offline function call tracing for doing audits whereas LLMs yield a zero profile approach resulting in a simplified software process.
8. Data Availability Statement
This work involves extensive experimentation with large language models (Claude 4 Sonnet, GPT-4, GPT-5) and other machine learning models, combined with static analysis and profiling techniques. We rely on established benchmark suites such as SPEC CPU 2017 and GAPS, and use a variety of profiling tools to establish ground truths for evaluating model predictions. We plan to submit our artifacts for the Artifact Evaluation process. While the artifacts are still being finalized, we intend to make available: (i) the source code for the benchmarks used; (ii) static analysis outputs derived from these codebases; (iii) model predictions; and (iv) evaluation scripts. These materials will enable reviewers and researchers to reproduce and build upon our results. Preliminary versions of the artifacts will be anonymized and linked as part of the artifact submission.